# Radiation forces and torques in optics and acoustics

### v1

@@ Acknowledgements.
We thank M. Nieto-Vesperinas and M. Petrov for fruitful discussions. This work was supported by the Australian Research Council (Grant No. DP210101292), the International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (contract FA520923C0023), EIC-Pathfinder-CHIRALFORCE (101046961) which is funded by Innovate UK Horizon Europe Guarantee (UKRI project 10045438), Ikerbasque (Basque Foundation of Science), Marie Skłodowska-Curie COFUND Programme of the European Commission (project HORIZON-MSCA-2022-COFUND-101126600-SmartBRAIN3), the International Research Agendas Programme (IRAP) of the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund and Teaming Horizon 2020 program of the European Commission [ENSEMBLE3 Project No. MAB/2020/14], and the project of the Ministry of Science and Higher Education (Poland) “Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program” [contract MEiN/2023/DIR/3797], Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2], the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).References
@@
- Alu and Engheta [2010] Alu, A, and N. Engheta (2010), “How does zero forward-scattering in magnetodielectric nanoparticles comply with the optical theorem?” p. 041590.
@@
- Barton et al. [1989] Barton, J P, D. R. Alexander, and S. A. Schaub (1989), “Theoretical determination of net radiation force and torque for a spherical particle illuminated by a focused laser beam,” J. Appl. Phys. 66, 4594–4602, arXiv:cs/9605103 .
@@
- Jiang et al. [2016] Jiang, Y, J. Chen, J. Ng, and Z. Lin (2016), “Decomposition of optical force into conservative and nonconservative components,” arXiv 1604.05138 .
@@
- Li et al. [2015] Li, P, Z. Mao, Z. Peng, L. Zhou, Y. Chen, P.-H. Huang, C. I. Truica, J. J. Drabick, W. S. El-Deiry, M. Dao, S. Suresh, and T. J. Huang (2015), “Acoustic separation of circulating tumor cells,” Proc. Natl. Acad. Sci. U.S.A. 112, 4970–4975.
@@
- Li et al. [2019b] Li, X, J. Chen, Z. Lin, and J. Ng (2019b), “Optical pulling at macroscopic distances,” Sci. Adv. 5, eaau7814.
@@
- Russell [2011] Russell, D (2011), “Acoustic Propulsion,” .
@@
- Silva et al. [2012a] Silva, G T, A. L. Baggio, J. H. Lopes, and F. G. Mitri (2012a), “Exact computations of the acoustic radiation force on a sphere using the translational addition theorem,” IEEE 62, 576–583, arXiv:1210.2116 .
@@
- Silva et al. [2012b] Silva, G T, T. P. Lobo, and F. G. Mitri (2012b), “Radiation torque produced by an arbitrary acoustic wave,” Europhys. Lett. 97, 54003.
@@
- Smagin et al. [2023] Smagin, M, I. Toftul, K. Y. Bliokh, and M. Petrov (2023), “Acoustic Lateral Recoil Force and Stable Lift of Anisotropic Particles,” arXiv 10.48550/arXiv.2310.06524, 2310.06524 .
@@
- Wang et al. [2024a] Wang, B, Z. Che, C. Cheng, C. Tong, L. Shi, Y. Shen, K. Y. Bliokh, and J. Zi (2024a), “Topological water-wave structures manipulating particles,” arXiv 10.48550/arXiv.2406.08515, 2406.08515 .
- Wang et al. [2024b] Wang, N, J. Ng, and G. P. Wang (2024b), “Morphology-independent general-purpose optical surface tractor beam,” Nat. Commun. 15, 6836.
@@
- Yang et al. [2023] Yang, C, D. Zhang, J. Zhao, W. Gao, W. Yuan, Y. Long, Y. Pan, H. Chen, F. Nori, K. Y. Bliokh, Z. Zhong, and J. Ren (2023), “Hybrid Spin and Anomalous Spin-Momentum Locking in Surface Elastic Waves,” Phys. Rev. Lett. 131, 136102.

### v3

@@ Acknowledgements.
We thank Etienne Brasselet, Xiao Li, Manuel Nieto-Vesperinas, and Mikhail Petrov for helpful remarks. This work was supported by the Australian Research Council (Grant No. DP210101292), the International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (contract FA520923C0023), EIC-Pathfinder-CHIRALFORCE (101046961) which is funded by Innovate UK Horizon Europe Guarantee (UKRI project 10045438), Ikerbasque (Basque Foundation of Science), Marie Skłodowska-Curie COFUND Programme of the European Commission (project HORIZON-MSCA-2022-COFUND-101126600-SmartBRAIN3), the International Research Agendas Programme (IRAP) of the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund and Teaming Horizon 2020 program of the European Commission [ENSEMBLE3 Project No. MAB/2020/14], and the project of the Ministry of Science and Higher Education (Poland) “Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program” [contract MEiN/2023/DIR/3797], Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2], the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).References
@@
- Aggarwal and Salapaka [2010] Aggarwal, T, and M. Salapaka (2010), “Real-time nonlinear correction of back-focal-plane detection in optical tweezers,” Rev. Sci. Instrum. 81 (12), 123105.
@@
- Allen [1966] Allen, P J (1966), “A radiation torque experiment,” Am. J. Phys. 34, 1185–1192.
@@
- Alu and Engheta [2010] Alu, A, and N. Engheta (2010), “How does zero forward-scattering in magnetodielectric nanoparticles comply with the optical theorem?” J. Nanophotonics 4, 041590.
@@
- Argun et al. [2020] Argun, A, T. Thalheim, S. Bo, F. Cichos, and G. Volpe (2020), “Enhanced force-field calibration via machine learning,” Appl. Phys. Rev. 7 (4), 10.1063/5.0019105.
@@
- Barton et al. [1989] Barton, J P, D. R. Alexander, and S. A. Schaub (1989), “Theoretical determination of net radiation force and torque for a spherical particle illuminated by a focused laser beam,” J. Appl. Phys. 66, 4594–4602.
@@
- Bliokh [2024] Bliokh, K Y (2024), “Momentum, spin, and orbital angular momentum of electromagnetic, acoustic, and water waves,” Contemp. Phys. 65, 219–238.
@@
- Bliokh et al. [2025] Bliokh, K Y, Z. Kuang, and S. Rotter (2025), “Dynamic and geometric shifts in wave scattering,” Rep. Prog. Phys. 88, 107901.
@@
- Bronte Ciriza et al. [2023] Bronte Ciriza, D, A. Magazzù, A. Callegari, G. Barbosa, A. A. R. Neves, M. A. Iatì, G. Volpe, and O. M. Maragò (2023), “Faster and More Accurate Geometrical-Optics Optical Force Calculation Using Neural Networks,” ACS Photonics 10 (1), 234–241.
@@
- Cameron et al. [2024] Cameron, R P, D. McArthur, A. M. Yao, N. Vogeley, and D. Wang (2024), “Robust chiral optical force via electric dipole interactions, inspired by a sea creature,” arXiv 2412.13206 .
@@
- Chen et al. [2020] Chen, H, H. Zheng, W. Lu amd S. Liu, J. Ng, and Z. Lin (2020), “Lateral Optical Force due to the Breaking of Electric-Magnetic Symmetry,” Phys. Rev. Lett. 125, 073901.
@@
- Ciarlo et al. [2024] Ciarlo, A, D. B. Ciriza, M. Selin, O. M. Maragò, A. Sasso, G. Pesce, G. Volpe, and M. Goksör (2024), “Deep learning for optical tweezers,” Nanophotonics 13 (17), 3017–3035.
@@
- Domina et al. [2025] Domina, K, P. Alonso-González, A. Bylinkin, M. Barra-Burillo, A. I. F. Tresguerres-Mata, F. Javier Alfaro-Mozaz, S. Vélez, F. Casanova, L. E. Hueso, R. Hillenbrand, K. Y. Bliokh, and A. Y. Nikitin (2025), “High-intensity wave vortices around subwavelength holes: From ocean tides to nanooptics,” Newton 1, 100060.
@@
- Falkovich et al. [2005] Falkovich, G, A. Weinberg, P. Denissenko, and S. Lukaschuk (2005), “A radiation torque experiment,” Nature 435, 1045–1046.
@@
- Fernandez-Corbaton et al. [2013] Fernandez-Corbaton, I, X. Zambrana-Puyalto, N. Tischler, X. Vidal, M. L. Juan, and G. Molina-Terriza (2013), “Electromagnetic duality symmetry and helicity conservation for the macroscopic maxwell’s equations,” Phys. Rev. Lett. 111, 060401.
- Ferretti et al. [2024] Ferretti, S, S. Marrara, D. Bronte Ciriza, A. Magazzù, A. Foti, P. G. Gucciardi, A. Musolino, L. Folco, V. Della Corte, A. Rotundi, R. Saija, A. Mandanici, O. M. Maragò, and M. G. Donato (2024), “Contactless manipulation and raman analysis of cometary analogs and micrometeorites by acoustic levitation,” Astrophys. J. 974, 287.
@@
- Guzatov and Klimov [2011] Guzatov, D V, and V. V. Klimov (2011), “Chiral particles in a circularly polarised light field: new effects and applications,” Quantum Electron. 41, 526.
@@
- Helgadottir et al. [2019] Helgadottir, S, A. Argun, and G. Volpe (2019), “Digital video microscopy enhanced by deep learning,” Optica 6 (4), 506–513.
@@
- Huan et al. [2024] Huan, Z, J. Zhou, Y. Xie, J. Xu, H. Wang, W. Ma, X. Li, W. Zhou, and T. Luo (2024), “Automated droplet manipulation enabled by a machine-vision-assisted acoustic tweezer,” Sens. Actuators, B 418, 136352.
@@
- Igoshin et al. [2025] Igoshin, V, A. Kokhanovskiy, and M. Petrov (2025), “Inverse design of Mie resonators with minimal backscattering,” Opt. Lett. 50 (5), 1735–1738.
@@
- Jiang et al. [2010] Jiang, Y, T. Narushima, and H. Okamoto (2010), “Nonlinear optical effects in trapping nanoparticles with femtosecond pulses,” Nat. Phys. 6, 1005–1009.
@@
- Jooss et al. [2022] Jooss, V M, J. S. Bolten, J. Huwyler, and D. Ahmed (2022), “In vivo acoustic manipulation of microparticles in zebrafish embryos,” Sci. Adv. 8, eabm2785.
@@
- Juan et al. [2009] Juan, M L, R. Gordon, Y. Pang, F. Eftekhari, and R. Quidant (2009), “Self-induced back-action optical trapping of dielectric nanoparticles,” Nat. Phys. 5, 915.
@@
- Kleshchenko et al. [2024] Kleshchenko, V, K. Albitskaya, and M. Petrov (2024), “Acoustic forces near elastic substrate,” Appl. Phys. Lett. 125, 212203.
- Kondo and Okubo [2022] Kondo, S, and K. Okubo (2022), “Improved mid-air acoustic tweezers using adaptive phase and amplitude control,” Jpn. J. Appl. Phys. 61 (SG), SG8004.
@@
- Lee and Chae [2024] Lee, Y, and E. Chae (2024), “Machine learning-enhanced optical tweezers for defect-free rearrangement,” Curr. Appl. Phys. 61, 150–159.
@@
- Lenton et al. [2020] Lenton, I C D, G. Volpe, A. B. Stilgoe, T. A. Nieminen, and H. Rubinsztein-Dunlop (2020), “Machine learning reveals complex behaviours in optically trapped particles,” Mach. Learn.: Sci. Technol. 1 (4), 045009.
@@
- Li et al. [2019b] Li, N, J. Cadusch, K. Crozier, and K. Crozier (2019b), “Algorithmic approach for designing plasmonic nanotweezers,” Opt. Lett. 44 (21), 5250–5253.
- Li et al. [2021a] Li, N, J. Cadusch, A. Liu, A. J. Barlow, A. Roberts, and K. B. Crozier (2021a), “Algorithm-Designed Plasmonic Nanotweezers: Quantitative Comparison by Theory, Cathodoluminescence, and Nanoparticle Trapping,” Adv. Opt. Mater. 9 (19), 2100758.
- Li et al. [2015a] Li, P, Z. Mao, Z. Peng, L. Zhou, Y. Chen, P.-H. Huang, C. I. Truica, J. J. Drabick, W. S. El-Deiry, M. Dao, S. Suresh, and T. J. Huang (2015a), “Acoustic separation of circulating tumor cells,” Proc. Natl. Acad. Sci. U.S.A. 112, 4970–4975.
@@
- Li et al. [2024] Li, X, Y. Cao, and J. Ng (2024), “Non-hermitian non-equipartition theory for trapped particles,” Nat. Commun. 15, 1963.
- Li et al. [2019c] Li, X, J. Chen, Z. Lin, and J. Ng (2019c), “Optical pulling at macroscopic distances,” Sci. Adv. 5, eaau7814.
- Li et al. [2021b] Li, X, Y. Liu, Z. Lin, J. Ng, and C. T. Chan (2021b), “Non-hermitian physics for optical manipulation uncovers inherent instability of large clusters,” Nat. Commun. 12, 6597.
- Li et al. [2015b] Li, Y, A. V. Maslov, N. I. Limberopoulos, A. M. Urbas, and V. N. Astratov (2015b), “Spectrally resolved resonant propulsion of dielectric microspheres,” Laser Photonics Rev. 9, 263–273.
- Li et al. [2013] Li, Y, O. V. Svitelskiy, A. V. Maslov, D. Carnegie, E. Rafailov, and V. N. Astratov (2013), “Giant resonant light forces in microspherical photonics,” Light Sci. Appl. 2, e64.
@@
- Lin et al. [2025] Lin, R, H.-S. Zhong, Y. Li, Z.-R. Zhao, L.-T. Zheng, T.-R. Hu, H.-M. Wu, Z. Wu, W.-J. Ma, Y. Gao, Y.-K. Zhu, Z.-F. Su, W.-L. Ouyang, Y.-C. Zhang, J. Rui, M.-C. Chen, C.-Y. Lu, and J.-W. Pan (2025), “AI-Enabled Parallel Assembly of Thousands of Defect-Free Neutral Atom Arrays,” Phys. Rev. Lett. 135 (6), 060602.
@@
- Liška et al. [2024] Liška, V, T. Zemánková, P. Jákl, M. Šiler, S. H. Simpson, P. Zemánek, and O. Brzobohatý (2024), “PT-like phase transition and limit cycle oscillations in non-reciprocally coupled optomechanical oscillators levitated in vacuum,” Nat. Phys. 20, 1622–1628.
@@
- Magallanes and Brasselet [2018] Magallanes, H, and E. Brasselet (2018), “Macroscopic direct observation of optical spin-dependent lateral forces and left-handed torques,” Nat. Photonics 12, 461.
@@
- Maslov [2014] Maslov, A V (2014), “Resonant Pulling of a Microparticle Using a Backward Surface Wave,” Phys. Rev. Lett. 112, 113903.
- Maslov [2020] Maslov, A V (2020), “Binding of Resonant Dielectric Particles to Metal Surfaces Using Plasmons,” Ann. Phys. 532, 2000129.
@@
- Medany et al. [2025] Medany, M, L. Piglia, L. Achenbach, S. K. Mukkavilli, and D. Ahmed (2025), “Model-based reinforcement learning for ultrasound-driven autonomous microrobots,” Nat. Mach. Intell. 7, 1076–1090.
@@
- Mirzaei-Ghormish et al. [2025] Mirzaei-Ghormish, S, D. Smalley, and R. Camacho (2025), “Nonlinear multistable potential traps,” Phys. Rev. A 111 (1), 013514.
@@
- Morrell et al. [2024] Morrell, M C, J. E. Lee, and D. G. Grier (2024), “Spectral holographic trapping: Creating dynamic force landscapes with polyphonic waves,” Phys. Rev. E 109, 044901.
@@
- Otte and Denz [2020] Otte, E, and C. Denz (2020), “Optical trapping gets structure: Structured light for advanced optical manipulation,” Appl. Phys. Rev. 7, 041308.
@@
- Patti et al. [2019] Patti, F, R. Saija, P. Denti, G. Pellegrini, P. Biagioni, M. A. Iatì, and O. M. Maragò (2019), “Chiral route to pulling optical forces and left-handed optical torques,” Sci. Rep. 9, 29.
@@
- Rakich et al. [2009] Rakich, P T, M. A. Popović, and Z. Wang (2009), “General treatment of optical forces and potentials in mechanically variable photonic systems,” Opt. Express 17, 18116–18135.
@@
- Raymond et al. [2020] Raymond, S J, D. J. Collins, R. O’Rorke, M. Tayebi, Y. Ai, and J. Williams (2020), “A deep learning approach for designed diffraction-based acoustic patterning in microchannels,” Sci. Rep. 10 (8745), 1–12.
@@
- Reisenbauer et al. [2024] Reisenbauer, M, H. Rudolph, L. Egyed, K. Hornberger, A. V. Zasedatelev, M. Abuzarli, B. A. Stickler, and U. Delić (2024), “Non-hermitian dynamics and non-reciprocity of optically coupled nanoparticles,” Nat. Phys. 20, 1629–1635.
- Ren et al. [2024] Ren, Z, X. Yan, K. Wen, H. Chen, E. Hajiyev, C. He, and G.-B. Jo (2024), “Creation of a tweezer array for cold atoms utilizing a generative neural network,” APL Quantum 1 (4), 10.1063/5.0219170.
@@
- Rieser et al. [2022] Rieser, J, M. A. Ciampini, H. Rudolph, N. Kiesel, K. Hornberger, B. A. Stickler, M. Aspelmeyer, and U. Delić (2022), “Tunable light-induced dipole-dipole interaction between optically levitated nanoparticles,” Science 377 (6609), 987–990.
- Rockstuhl and Herzig [2004] Rockstuhl, C, and H. P. Herzig (2004), “Wavelength-dependent optical force on elliptical silver cylinders at plasmon resonance,” Opt. Lett. 29, 2181–2183.
@@
- Rudolph et al. [2022] Rudolph, H, U. Delić, M. Aspelmeyer, K. Hornberger, and B. A. Stickler (2022), “Force-gradient sensing and entanglement via feedback cooling of interacting nanoparticles,” Phys. Rev. Lett. 129, 193602.
@@
- Russell [2011] Russell, D (2011), “Acoustic Propulsion,” youtube.com/watch?v=je7eLZS6GG0 .
- Russell [2020] Russell, D (2020), “Acoustic Radiometer,” youtube.com/watch?v=KGLFJzFt-G4 .
- Russell [2025] Russell, D A (2025), “Acoustic rockets: A modern demonstration of the 1878 acoustic repulsion apparatus by Dvorak and Mayer,” Proc. Mtgs. Acoust. 42, 025003.
@@
- Schäfer et al. [2021] Schäfer, J, H. Rudolph, K. Hornberger, and B. A. Stickler (2021), “Cooling nanorotors by elliptic coherent scattering,” Phys. Rev. Lett. 126, 163603.
@@
- Schrage et al. [2023] Schrage, M, M. Medany, and D. Ahmed (2023), “Ultrasound Microrobots with Reinforcement Learning,” Adv. Mater. Technol. 8 (10), 2201702.
@@
- Shang et al. [2013] Shang, Q-C, Z.-S. Wu, T. Qu, Z.-J. Li, L. Bai, and L. Gong (2013), “Analysis of the radiation force and torque exerted on a chiral sphere by a gaussian beam,” Opt. Express 21, 8677–8688.
@@
- Shen and Liu [2025] Shen, Z, and N. Liu (2025), “Optical Tweezers with Optical Vortex Based on Deep Learning,” ACS Photonics 12 (4), 2212–2218.
@@
- Shi et al. [2023] Shi, Y, X. Xu, M. Nieto-Vesperinas, Q. Song, A. Q. Liu, G. Cipparrone, Z. Su, B. Yao, Z. Wang, C.-W. Qiu, and X. Cheng (2023), “Advances in light transverse momenta and optical lateral forces,” Adv. Opt. Photon. 15, 835–906.
@@
- Silva et al. [2015] Silva, G T, A. Baggio, J. H. Lopes, and F. G. Mitri (2015), “Computing the acoustic radiation force exerted on a sphere using the translational addition theorem,” IEEE Trans. Ultrason. Ferroelectr. Freq. Control 62, 576–583.
@@
- Silva et al. [2012] Silva, G T, T. P. Lobo, and F. G. Mitri (2012), “Radiation torque produced by an arbitrary acoustic wave,” Europhys. Lett. 97, 54003.
@@
- Sitters et al. [2015] Sitters, G, D. Kamsma, G. Thalhammer, M. Ritsch-Marte, E. J. G. Peterman, and G. J. L. Wuite (2015), “Acoustic force spectroscopy,” Nat. Methods 12, 47–50.
@@
- Smagin et al. [2024] Smagin, M, I. Toftul, K. Y. Bliokh, and M. Petrov (2024), “Acoustic lateral recoil force and stable lift of anisotropic particles,” Phys. Rev. Appl. 22, 064041.
@@
- Stout and Webb [1997] Stout, A L, and W. W. Webb (1997), “Optical force microscopy,” (Academic Press) pp. 99–116.
@@
- Toftul et al. [2025a] Toftul, I, Y. S. Kivshar, and M. Lapine (2025a), “Acoustic angular sorting of resonant subwavelength particles,” Phys. Rev. Applied 24, 034019.
- Toftul et al. [2020] Toftul, I, D. Kornovan, and M. Petrov (2020), “Self-Trapped Nanoparticle Binding via Waveguide Mode,” ACS Photonics 7, 114–119.
- Toftul et al. [2025b] Toftul, I, L. Mao, S. Balendhran, M. Taha, Y. Kivshar, and S. Kruk (2025b), “Hopping and orbiting of Mie-resonant nanoparticles in optical tweezers,” arXiv:2507.10058 10.48550/arXiv.2507.10058.
- Toftul et al. [2025c] Toftul, I, M. Petrov, R. Quidant, and Y. Kivshar (2025c), “Optical Supertorque Induced by Mie-Resonant Modes,” ACS Photonics 12, 3447–3456.
@@
- Wang et al. [2025] Wang, B, Z. Che, C. Cheng, C. Tong, L. Shi, Y. Shen, K. Y. Bliokh, and J. Zi (2025), “Topological water-wave structures manipulating particles,” Nature 638, 394––400.
- Wang et al. [2024] Wang, N, J. Ng, and G. P. Wang (2024), “Morphology-independent general-purpose optical surface tractor beam,” Nat. Commun. 15, 6836.
@@
- Wang et al. [2023] Wang, S, X. Wang, F. You, and H. Xiao (2023), “A Deep Learning Based Method for Generating Holographic Acoustic Fields from Phased Transducer Arrays,” in Artificial Neural Networks and Machine Learning – ICANN 2023, pp. 13–24.
@@
- Wu and Liao [2024] Wu, M, and W. Liao (2024), “Machine Learning-Empowered Real-Time Acoustic Trapping: An Enabling Technique for Increasing MRI-Guided Microbubble Accumulation,” Sensors 24 (19), 6342.
@@
- Wu et al. [2025] Wu, Y-J, J.-H. Zhuang, P.-P. Yu, Y.-F. Liu, Z.-Q. Wang, Y.-M. Li, C.-W. Qiu, and L. Gong (2025), “Time-varying 3d optical torque via a single beam,” Nat. Commun. 16, 593.
@@
- Yang et al. [2023a] Yang, C, D. Zhang, J. Zhao, W. Gao, W. Yuan, Y. Long, Y. Pan, H. Chen, F. Nori, K. Y. Bliokh, Z. Zhong, and J. Ren (2023a), “Hybrid Spin and Anomalous Spin-Momentum Locking in Surface Elastic Waves,” Phys. Rev. Lett. 131, 136102.
- Yang et al. [2025] Yang, M, Y. Shi, Q. Song, Z. Wei, X. Dun, Z. Wang, Z. Wang, C.-W. Qiu, H. Zhang, and X. Cheng (2025), “Optical sorting: past, present and future,” Light Sci. Appl. 14, 103.
@@
- Yang et al. [2023b] Yang, Y, Y. Yang, D. Liu, Y. Wang, M. Lu, Q. Zhang, J. Huang, Y. Li, T. Ma, F. Yan, and H. Zheng (2023b), “In-vivo programmable acoustic manipulation of genetically engineered bacteria,” Nat. Commun. 14, 3297.
@@
- Yu et al. [2019] Yu, X, Y. Jiang, H. Chen, S. Liu, and Z. Lin (2019), “Approach to fully decomposing an optical force into conservative and nonconservative components,” Phys. Rev. A 100, 033821.
@@
- Zhang et al. [2018] Zhang, Y, J. Shen, C. Min, Y. Jin, Y. Jiang, J. Liu, S. Zhu, Y. Sheng, A. V. Zayats, and X. Yuan (2018), “Nonlinearity-Induced Multiplexed Optical Trapping and Manipulation with Femtosecond Vector Beams,” Nano Lett. 18, 5538–5543.
@@
- Zheng et al. [2020a] Zheng, H, X. Li, Y. Jiang, J. Ng, Z. Lin, and H. Chen (2020a), “General formulations for computing the optical gradient and scattering forces on a spherical chiral particle immersed in generic monochromatic optical fields,” Phys. Rev. A 101, 053830.
- Zheng et al. [2020b] Zheng, H, X. Li, J. Ng, H. Chen, and Z. Lin (2020b), “Tailoring the gradient and scattering forces for longitudinal sorting of generic-size chiral particles,” Opt. Lett. 45, 4515–4518.
@@
- Zhu et al. [2023] Zhu, L, Y. Tai, H. Li, H. Hu, X. Li, Y. Cai, and Y. Shen (2023), “Multidimensional optical tweezers synthetized by rigid-body emulated structured light,” Photonics Research 11, 1524–1534.